Reduce the following fractions to the simplest form: 
12 / 52 

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

12 / 52 = (3 × 4) / (13 × 4)
= 3/13

Thanks
Shraddha Singh
one year ago
Dear Student,
 
The simplest form of the fraction 12/52 is
12/52 = 12÷4/52÷4 = 3/13
 
Hope it helps!
Thank you